The surface area of a cube is 150 sq. cm. What is the length of its diagonal in cm?
Answer by Alan3354(21584)

The surface area of a cube is 150 sq. cm. What is the length of its diagonal in cm?
s = length of side
A cube has 6 faces, each has an area of s^2
6s^2 = 150
s^2 = 25
s = 5
-----------
The diagonal across a face is s*sqrt(2) = 5sqrt(2). Any question about that, email via the thank you note.
-----------------
d^2 = 5^2 + (5sqrt(2))^2 = 25 + 50
d^2 = 75
d = 5*sqrt(3)